An adult female who weighs 164 lb consumes on average 2.830×103 Cal/day. If she wants to lose 30.0 lb in 357 days , what should her daily caloric intake (Calories per day) be to achieve this weight loss?
Answer #1

For the loss of 1lb calories that has to be burnt out is 3500 cal

hence calories required to burn 30 lb = 30 * 3500 = 105000 cal

calories that needed to be burnt per day = 105000 / 357 = 294.11 cal

now her daily intake would be 2.830 * 103 - 294.11 = 2535.89 cal / day